OpenStudy (anonymous):

I don't understand why this problem is considered unsolvable: (-4)^1/2

OpenStudy (accessdenied):

What were you considering for a solution?

OpenStudy (anonymous):

It says the answer is suppose to be unsolvable but I don't know why

OpenStudy (accessdenied):

Were you considering that (-4)^(1/2) is the same as the square root of -4 ? \( (-4)^{1/2} = \sqrt{-4} \)

OpenStudy (anonymous):

Um, idk, I guess yes I was. I'm teaching this stuff to myself and the text didn't explain

OpenStudy (anonymous):

So, because its not a sq. root but rather a simple negative number...

OpenStudy (accessdenied):

You said you don't know why this is unsolvable. So I was wondering if you had an idea for how it might be solved, so we can go over why it might not work!

OpenStudy (accessdenied):

Let's call our expression, y. y = (-4)^(1/2) Now we can square both sides. y^2 = - 4 So, think about what you could put in for y whose square is negative four. If we put in a positive number, its square is also a positive number. 2^2 = +4. If we put in a negative number, it is ALSO a positive number because we multiply -2 by -2. The negatives cancel! (-2)^2 = (-2)(-2) = 4. This is where real number solutions do not exist. However, if you go further you might learn about complex numbers, whose basis is the "number" i = sqrt(-1). But in our case, unsolvable means "not a real number".